\begin{thebibliography}{99}

%\cite{Ortin:2021ade}
\bibitem{Ortin:2021ade}
T.~Ort\'{\i}n,
``Komar integrals for theories of higher order
in the Riemann curvature and black-hole chemistry,''
JHEP \textbf{08} (2021), 023
\doi{10.1007/JHEP08(2021)023}
[\arxiv{2104.10717} [gr-qc]].

%\cite{Mitsios:2021zrn}
\bibitem{Mitsios:2021zrn}
D.~Mitsios, T.~Ort\'{\i}n and D.~Pere\~niguez,
``Komar integral and Smarr formula for axion-dilaton
black holes versus S duality,''
JHEP \textbf{08} (2021), 019
\doi{10.1007/JHEP08(2021)019}
[\arxiv{2106.07495} [hep-th]].


%\cite{Meessen:2022hcg}
\bibitem{Meessen:2022hcg}
P.~Meessen, D.~Mitsios and T.~Ort\'{\i}n,
``Black hole chemistry, the cosmological constant and the embedding tensor,''
JHEP \textbf{12} (2022), 155
\doi{10.1007/JHEP12(2022)155}
[\arxiv{2203.13588} [hep-th]].

%\cite{Ortin:2022uxa}
\bibitem{Ortin:2022uxa}
T.~Ort\'{\i}n and D.~Pere\~niguez,
``Magnetic charges and Wald entropy,''
JHEP \textbf{11} (2022), 081
\doi{10.1007/JHEP11(2022)081}
[\arxiv{2207.12008} [hep-th]].

%\cite{Ballesteros:2023iqb}
\bibitem{Ballesteros:2023iqb}
R.~Ballesteros, C.~G\'omez-Fayr\'en, T.~Ort\'{\i}n and M.~Zatti,
``On scalar charges and black hole thermodynamics,''
JHEP \textbf{05} (2023), 158
\doi{10.1007/JHEP05(2023)158}
[\arxiv{2302.11630} [hep-th]].

%\cite{Gomez-Fayren:2023wxk}
\bibitem{Gomez-Fayren:2023wxk}
C.~Gomez-Fayren, P.~Meessen, T.~Ortin and M.~Zatti,
``Wald entropy in Kaluza-Klein black holes,''
JHEP \textbf{08} (2023), 039
\doi{10.1007/JHEP08(2023)039}
[\arxiv{2305.01742} [hep-th]].

%\cite{Bandos:2023zbs}
\bibitem{Bandos:2023zbs}
I.~Bandos and T.~Ort\'{\i}n,
``Noether-Wald charge in supergravity: the fermionic contribution,''
JHEP \textbf{12} (2023), 095
\doi{10.1007/JHEP12(2023)095}
[\arxiv{2305.10617} [hep-th]].

%\cite{Ballesteros:2023muf}
\bibitem{Ballesteros:2023muf}
R.~Ballesteros and T.~Ort\'{\i}n,
``Hairy black holes, scalar charges and extended thermodynamics,''
Class. Quant. Grav. \textbf{41} (2024) no.5, 055007
\doi{10.1088/1361-6382/ad210a}
[\arxiv{2308.04994} [gr-qc]].

%\cite{Ballesteros:2024prz}
\bibitem{Ballesteros:2024prz}
R.~Ballesteros and T.~Ort\'{\i}n,
``Generalized Komar charges and Smarr formulas
for black holes and boson stars,''
[\arxiv{2409.08268} [gr-qc]].

%\cite{Bandos:2024pns}
\bibitem{Bandos:2024pns}
I.~Bandos, P.~Meessen and T.~Ort\'\i{}n,
``Noether-Wald and Komar charges in supergravity, fermions, and Killing supervectors in superspace,''
J. Phys. Conf. Ser. \textbf{2912} (2024) no.1, 012007
\doi{10.1088/1742-6596/2912/1/012007}
[\arxiv{2411.01020} [hep-th]].

%\cite{Ortin:2024emt}
\bibitem{Ortin:2024emt}
T.~Ort\'\i{}n and M.~Zatti,
``On the thermodynamics of the black holes of the Cano-Ruip\'erez 4-dimensional string effective action,''
[\arxiv{2411.10417} [hep-th]].

%\cite{Bandos:2024rit}
\bibitem{Bandos:2024rit}
I.~Bandos, P.~Meessen and T.~Ort\'\i{}n,
``Komar charge of N = 2 supergravity and its superspace generalization,''
[\arxiv{2412.18510} [hep-th]].

%\cite{Komar:1958wp}
\bibitem{Komar:1958wp}
A.~Komar,
``Covariant conservation laws in general relativity,''
Phys. Rev. \textbf{113} (1959), 934-936
\doi{10.1103/PhysRev.113.934}

%\cite{Elgood:2020svt}
\bibitem{Elgood:2020svt}
Z.~Elgood, P.~Meessen and T.~Ort\'\i{}n,
``The first law of black hole mechanics in
the Einstein-Maxwell theory revisited,''
JHEP \textbf{09} (2020), 026
\doi{10.1007/JHEP09(2020)026}
[\arxiv{2006.02792} [hep-th]].

%\cite{Elgood:2020mdx}
\bibitem{Elgood:2020mdx}
Z.~Elgood, D.~Mitsios, T.~Ort\'{\i}n and D.~Pere\~niguez,
``The first law of heterotic stringy black hole mechanics
at zeroth order in \ensuremath{\alpha}',''
JHEP \textbf{07} (2021), 007
\doi{10.1007/JHEP07(2021)007}
[\arxiv{2012.13323} [hep-th]].

%\cite{Elgood:2020nls}
\bibitem{Elgood:2020nls}
Z.~Elgood, T.~Ort\'{\i}n and D.~Pere\~niguez,
``The first law and Wald entropy formula of
heterotic stringy black holes at first order in $\alpha'$,''
JHEP \textbf{05} (2021), 110
\doi{10.1007/JHEP05(2021)110}
[\arxiv{2012.14892} [hep-th]].

%\cite{Ortin:2015hya}
\bibitem{Ortin:2015hya}
T.~Ort\'{\i}n,
``Gravity and Strings'', 2nd edition, 
Cambridge University Press, 2015.


%\cite{Barnich:2001jy}
\bibitem{Barnich:2001jy}
G.~Barnich and F.~Brandt,
``Covariant theory of asymptotic symmetries,
conservation laws and central charges,''
Nucl. Phys. B \textbf{633} (2002), 3-82
\doi{10.1016/S0550-3213(02)00251-1}
[\hepth{0111246} [hep-th]].

%\cite{Kastor:2008xb}
\bibitem{Kastor:2008xb}
D.~Kastor,
``Komar Integrals in Higher (and Lower) Derivative Gravity,''
Class. Quant. Grav. \textbf{25} (2008), 175007
\doi{10.1088/0264-9381/25/17/175007}
[\arxiv{0804.1832} [hep-th]].

%\cite{Kastor:2010gq}
\bibitem{Kastor:2010gq}
D.~Kastor, S.~Ray and J.~Traschen,
``Smarr Formula and an Extended First Law for Lovelock Gravity,''
Class. Quant. Grav. \textbf{27} (2010), 235014
\doi{10.1088/0264-9381/27/23/235014}
[\arxiv{1005.5053} [hep-th]].

%\cite{Liberati:2015xcp}
\bibitem{Liberati:2015xcp}
S.~Liberati and C.~Pacilio,
``Smarr Formula for Lovelock Black Holes: a Lagrangian approach,''
Phys. Rev. D \textbf{93} (2016) no.8, 084044
\doi{10.1103/PhysRevD.93.084044}
[\arxiv{1511.05446} [gr-qc]].

%\cite{Adami:2024gdx}
\bibitem{Adami:2024gdx}
  H.~Adami, M.~Golshani, M.~M.~Sheikh-Jabbari,
  V.~Taghiloo and M.~H.~Vahidinia,
``Covariant Phase Space Formalism for Fluctuating Boundaries,''
[\arxiv{2407.03259} [hep-th]].


\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage{graphicx}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{textcomp}
\usepackage[sc,osf]{mathpazo}
\usepackage{a4wide}  
\usepackage{latexsym,amsthm,amsfonts,amsmath,mathrsfs,amssymb}
\usepackage{dsfont}
\usepackage{accents}
\usepackage{xcolor}
\usepackage[nosort]{cite}
\usepackage{booktabs} 
\usepackage[unicode,implicit]{hyperref}
\hypersetup{%
  pdftitle    = {A note on the Noether--Wald and generalized Komar charges}
  pdfkeywords = {symmetry, conserved charges, Noether, Komar},
  pdfauthor   = {Tomas Ortin and Matteo Zatti},
  plainpages  = true,
  colorlinks  = true,
  citecolor   = blue,
  urlcolor    = red,
  linkcolor   = black
}
\newcommand{\hepth}[1]{{\tt
\href{http://www.arXiv.org/abs/hep-th/#1}{hep-th/#1}}}

\newcommand{\arxiv}[1]{{\tt arXiv:\href{http://www.arXiv.org/abs/#1}{#1}}}
\newcommand{\doi}[1]{{\tt DOI:\href{http://dx.doi.org/#1}{#1}}}

\allowdisplaybreaks
\begin{document}

\begin{flushright}
\small
IFT-UAM/CSIC-24-161\\
MPP-2024-222 \\
November 15\textsuperscript{th}, 2024\\
\normalsize
\end{flushright}

\vspace{1cm}

\begin{center}

  {\Large {\bf A note on the Noether--Wald\\[.5cm] and generalized Komar charges}}
 
\vspace{2cm}

\renewcommand{\thefootnote}{\alph{footnote}}

{\sl\large Tom\'{a}s Ort\'{\i}n}$^{1,}$\footnote{Email:
	{\tt tomas.ortin[at]csic.es}}
{\sl\large and Matteo Zatti}$^{1,2,}$\footnote{Email:
	{\tt zatti[at]mpp.mpg.de}}

\setcounter{footnote}{0}
\renewcommand{\thefootnote}{\arabic{footnote}}
\vspace{2cm}

${}^{1}${\it Instituto de F\'{\i}sica Te\'orica UAM/CSIC\\
	C/ Nicol\'as Cabrera, 13--15,  C.U.~Cantoblanco, E-28049 Madrid, Spain}\\

\vspace{0.5cm}

${}^{2}${\it Max Planck Institut f\"ur Physik \\
	Boltzmannstrasse 8, 85748 Garching, Germany}\\

\vspace{2cm}

{\bf Abstract}
\end{center}
\begin{quotation}
  {\small \noindent We propose a simple algorithm to compute the generalized Komar
    charge of any exactly gauge- and diffeomorphism-invariant theory.}
\end{quotation}

\newpage
\pagestyle{plain}

\noindent Conserved charges can be used to characterize the states and the evolution of physical systems. In this short note we describe how to compute the generalized Komar charge of any exactly gauge- and diffeomorphism-invariant theory. Our goal is to provide a simple algorithm which can be easily applied to spacetime configurations with Killing isometries (cf. Eq.~(\ref{eq:prescription})). This is particularly useful to determine efficiently fundamental identities in black hole thermodynamics and to prove non-existence theorems for gravitational solitons and boson stars (see for instance Refs.~\cite{Ortin:2021ade,Mitsios:2021zrn,Meessen:2022hcg,Ortin:2022uxa,Ballesteros:2023iqb, Gomez-Fayren:2023wxk, Bandos:2023zbs,Ballesteros:2023muf,Ballesteros:2024prz, Bandos:2024pns,Ortin:2024emt, Bandos:2024rit}).

In $d$-dimensional General Relativity, for each of the Killing vectors $k$ of
the metric of a given vacuum solution one can construct a $(d-2)$-form charge
$\mathbf{K}[k]$, the so-called \textit{Komar charge} \cite{Komar:1958wp},
which is closed on-shell\footnote{We indicate with $\doteq$ those identities
which are only satisfied on-shell.}
\begin{equation}
  d\mathbf{K}[k]
  \doteq
  0\,.
\end{equation}
In asymptotically-flat spacetimes, the integral of this charge at spatial
infinity ($S^{d-2}_{\infty}$) gives, up to normalization, the value of the
conserved charge of the spacetime corresponding to the Killing vector: total
mass/energy if $k$ is a Killing vector that generates time translations, etc.
In General Relativity, the Komar charge coincides with the Noether--Wald
charge associated with the invariance under diffeomorphisms generated by
vector fields $\xi$, $\mathbf{Q}[\xi]$, evaluated over the Killing vector $k$,
$\mathbf{Q}[k]$. This fact suggests that, in more general theories (with
matter or with terms of higher order in the curvature) the Komar charge may
also be given by the Noether--Wald charge evaluated on $k$, $\mathbf{Q}[k]$,
since $\mathbf{Q}[\xi]$ can be constructed in any theory invariant under
diffeomorphisms. This naive expectation turns out to be false in general. In order to understand why and what has to be done to construct an on-shell-closed 2-form (a \textit{generalized Komar charge}), it is convenient to review the
algorithm that leads to $\mathbf{Q}[\xi]$.

Let us consider a theory of gravity, described by the Vierbein $e^{a}$,
coupled to a number of matter fields denoted generically by $\varphi$, whose
dynamics is dictated by the action $S[e,\varphi]$.  Under a generic
infinitesimal variation of the fields
\begin{equation}
  \label{eq:generalvariation}
  \delta S[e,\varphi]
  =
  \int \left\{\mathbf{E}_{a}\wedge \delta e^{a}
    +\mathbf{E}_{\varphi}\wedge \delta\varphi
  +d\mathbf{\Theta}(e,\varphi,\delta e,\delta \varphi)\right\}\,,
\end{equation}
where, by definition, $\mathbf{E}_{a}$ are the Einstein equations,
$\mathbf{E}_{\varphi}$ are the equations of motion of the matter fields and
$\mathbf{\Theta}(e,\varphi,\delta e,\delta \varphi)$ is the \textit{symplectic
  prepotential}. If the theory is exactly invariant under diffeomorphisms and any other kind of gauge transformations,\footnote{We exclude Chern--Simons and any other terms
  which are invariant up to total derivatives we do not want to deal with. We
  must take into account that isometries induce gauge transformations, as
  explained in Refs.~\cite{Elgood:2020svt,Elgood:2020mdx,Elgood:2020nls}.}
\begin{equation}
  \label{eq:deltaxiS1}
  \delta_{\xi}S[e,\varphi]
  =
  -\int d\imath_{\xi}\mathbf{L}\,,
\end{equation}
where $\imath_{\xi}\mathbf{L}$ indicates the interior product of the vector
field $\xi$ with the $d$-form $\mathbf{L}$.\footnote{We are using
  differential-form notation. the rest of our conventions and notation can be
  found in Refs.~\cite{Elgood:2020svt, Ortin:2015hya}.}
On the other hand, if we use, instead, the general infinitesimal variation
Eq.~(\ref{eq:generalvariation}) particularized for infinitesimal
diffeomorphisms, we get
\begin{equation}
  \label{eq:deltaxiS2}
  \delta_{\xi} S[e,\varphi]
  =
  \int \left\{\mathbf{E}_{a}\wedge \delta_{\xi} e^{a}
    +\mathbf{E}_{\varphi}\wedge \delta_{\xi}\varphi
    +d\mathbf{\Theta}(e,\varphi,\delta_{\xi} e,\delta_{\xi} \varphi)\right\}\,.
\end{equation}
Combining equations \eqref{eq:deltaxiS1} and \eqref{eq:deltaxiS2} we arrive at the identity
\begin{equation} 
	\int \left\{\mathbf{E}_{a}\wedge \delta_{\xi} e^{a}
	+\mathbf{E}_{\varphi}\wedge \delta_{\xi}\varphi
	+d\left[\mathbf{\Theta}(e,\varphi,\delta_{\xi} e,\delta_{\xi} \varphi)
	+\imath_{\xi}\mathbf{L}\right]\right\}
	=
	0\,,
\end{equation}
which is satisfied off-shell for any $\xi$ and, more importantly, independently of the integration domain, which leads to
\begin{equation} \label{eq:step}
	\mathbf{E}_{a}\wedge \delta_{\xi} e^{a}
	+\mathbf{E}_{\varphi}\wedge \delta_{\xi}\varphi
	=
	d\left[-\mathbf{\Theta}(e,\varphi,\delta_{\xi} e,\delta_{\xi} \varphi)
	-\imath_{\xi}\mathbf{L}\right]\,.
\end{equation}
Equation \eqref{eq:step} tells us that the l.h.s.~is a total derivative. Such total derivative can be independently evaluated integrating by parts the l.h.s.~of \eqref{eq:step}. One must integrate all the terms involving derivatives of the gauge parameters until obtaining an expression that contains terms linear in the gauge parameters (without their derivatives) plus a total derivative
\begin{equation} \label{eq:step2}
		\mathbf{E}_{a}\wedge \delta_{\xi} e^{a}
		+\mathbf{E}_{\varphi}\wedge \delta_{\xi}\varphi
		=
		\mathcal{N}_{a}\xi^{a} +dN\,. 
\end{equation}
Comparing \eqref{eq:step} and \eqref{eq:step2} we conclude that the terms linear in the gauge parameters must vanish\footnote{Notice that $\mathcal{N}_a = 0$ are nothing but the  $\xi$-independent Noether identities. Our derivation is basically the derivation of the second Noether theorem.} and only the
total derivative $dN$ survives
\begin{equation}
\label{eq:Ndef}
\mathbf{E}_{a}\wedge \delta_{\xi} e^{a} +\mathbf{E}_{\varphi}\wedge
\delta_{\xi}\varphi
=
d N\,,
\end{equation}
where $N$ is a $(d-1)$-form that vanishes on-shell. Then, we can write
\begin{subequations}
  \label{eq:deltaxiS3}
  \begin{align}
  \delta_{\xi} S[e,\varphi]
  & =
    \int d\mathbf{\Theta}'\,,
    \\
    & \nonumber \\
    \label{eq:Thetaprimedef}
  \mathbf{\Theta}'
  & \equiv
\mathbf{\Theta}(e,\varphi,\delta_{\xi} e,\delta_{\xi} \varphi)+N\,,  
  \end{align}
\end{subequations}
and, comparing again this expression with Eq.~(\ref{eq:deltaxiS1}) one
concludes that the $(d-1)$-form
\begin{equation}
  \label{eq:Jxi}
  \mathbf{J}[\xi]
  \equiv
  \mathbf{\Theta}'+\imath_{\xi}\mathbf{L}\,,
\end{equation}
is closed off-shell for any vector field $\xi$
\begin{equation}
  d\mathbf{J}[\xi]
  =
  0\,.
\end{equation}
This implies the local existence of Noether--Wald $(d-2)$-form $\mathbf{Q}[k]$
defined, up to total derivatives, by
\begin{equation}
  d\mathbf{Q}[\xi]
  =
  \mathbf{J}[\xi]\,.
\end{equation}
Let us examine the right-hand side of this equation using the definition of
$\mathbf{J}[\xi]$, Eqs.~(\ref{eq:Jxi}), and that of $\mathbf{\Theta}'$,
Eq.~(\ref{eq:Thetaprimedef})
\begin{equation}
  d\mathbf{Q}[\xi]
  =
  \mathbf{\Theta}(e,\varphi,\delta_{\xi} e,\delta_{\xi} \varphi)+N
  +\imath_{\xi}\mathbf{L}\,.
\end{equation}
The second term in the right-hand side vanishes on-shell while the first
vanishes if we find \textit{Killing} (or \textit{reducibility}
\cite{Barnich:2001jy}) parameters\footnote{In general, the transformations
  $\delta_{\xi}$ depend on the vector field $\xi$ and other gauge
  parameters.}
$k$ such that
\begin{equation}
  \delta_{k} e^{a}
  =
  \delta_{k} \varphi
  =
  0\,,
\end{equation}
because the symplectic prepotential is linear on $\delta_{\xi} e^{a}$ and
$\delta_{\xi} \varphi$. The third term only vanishes in some simple theories
like General Relativity with no matter or with free, massless scalars, and, in
general, we have to deal with the equation
\begin{equation}
  \label{eq:OsdQk}
  d\left(\mathcal{O}_{s}\mathbf{Q}[k]\right)
  \doteq
  \mathcal{O}_{s}\imath_{k}\mathbf{L}\,,
\end{equation}
where we have introduced the on-shell-setting operator $\mathcal{O}_{s}$ that
evaluates the expression on its right over the solution $s$. Smarr formulas
can be computed integrating this identity over hypersurfaces with boundary on
the horizon and spatial infinity
\cite{Kastor:2008xb,Kastor:2010gq,Liberati:2015xcp,Adami:2024gdx}. In Refs.~\cite{Ortin:2021ade,Mitsios:2021zrn} it was argued that the
right-hand side of Eq.~(\ref{eq:OsdQk}) is always a total derivative 
\begin{equation}
  \mathcal{O}_{s}\imath_{k}\mathbf{L}
  \equiv
  d\omega_{k}\,,
\end{equation}
and we can define the \textit{generalized Komar $(d-2)$-form charge}
\begin{equation}
  \mathbf{K}[k]
  \equiv
  -\mathcal{O}_{s}\mathbf{Q}[k]
  +\omega_{k}\,,
\end{equation}
which is closed on-shell 
\begin{equation}
  d \mathbf{K}[k]
  \doteq
  0\,.
\end{equation}

At first sight, this construction may give a trivial $\mathbf{K}[k]$, but the
explicit calculations performed in
Refs.~\cite{Ortin:2021ade,Mitsios:2021zrn,Meessen:2022hcg,Ortin:2022uxa,Ballesteros:2023iqb, Gomez-Fayren:2023wxk, Bandos:2023zbs,Ballesteros:2023muf,Ballesteros:2024prz, Bandos:2024pns,Ortin:2024emt, Bandos:2024rit}
proof otherwise. However, a closer inspection of the way in which those
explicit calculations have been performed shows that what was computed in
those references is, actually, and more precisely,
\begin{equation}
  \label{eq:omegakdef2}
  \imath_{k}\mathcal{O}_{s}\mathbf{L}
  \equiv
  d\omega_{k}\,,
\end{equation}
where $\mathcal{O}_{s}\mathbf{L}$ is typically obtained from the trace of the
Einstein equations. The difference between this definition of $\omega_{k}$ and the former is a
total derivative that vanished on-shell
\begin{equation}
  \imath_{k}\mathcal{O}_{s}\mathbf{L}
  -\mathcal{O}_{s}\imath_{k}\mathbf{L}
  =
  d\left(\omega_{k}-\mathbf{Q}[k]\right)
  \doteq
  0\,,
\end{equation}
and it is, precisely, the total derivative of the generalized Komar charge. Thus, we arrive at the following prescription:
\begin{equation}
  \label{eq:prescription}
  d\mathbf{K}[k]
  =
  [\imath_{k},\mathcal{O}_{s}]\mathbf{L}\,.
\end{equation}

Let us see how the prescription works in the simple example of the
Einstein--Maxwell theory in $d$ dimensions. Its action is
\begin{equation}
\label{eq:EMaction2}
S[e^{a},A]
=
\frac{(-1)^{d-1}}{16\pi G_{N}^{(d)}}
\int 
\left[
\star (e^{a}\wedge e^{b}) \wedge R_{ab}  -\tfrac{1}{2}F\wedge \star F
\right]
\equiv
\int \mathbf{L}\,,
\end{equation}
and its equations of motion and symplectic prepotential, defined by
\begin{equation}
    \delta S
     =
     \int 
     \left\{
       \mathbf{E}_{a}\wedge \delta e^{a} +\mathbf{E}\wedge \delta A
       +d\mathbf{\Theta}(e,A,\delta e,\delta A)
    \right\}\,,
\end{equation}
are given by
\begin{subequations}
  \begin{align}
    \label{eq:Eb}
  \mathbf{E}_{a}
  & =
    \imath_{a}\star (e^{c}\wedge e^{d})\wedge R_{cd}
    +\tfrac{1}{2}\left(\imath_{a}F\wedge \star F-F\wedge \imath_{a}\star F\right)\,,
\\
    & \nonumber \\
    \label{eq:E}
    \mathbf{E}
    & =
      -d\star F\,,
    \\
    & \nonumber \\
      \mathbf{\Theta}(e,A,\delta e,\delta A)
  & =
  -\star (e^{a}\wedge e^{b})\wedge \delta \omega_{ab}
    +\star F \wedge \delta A\,,
  \end{align}
\end{subequations}
where $\imath_{c}$ stands for $\imath_{e_{c}}$, where
$e_{c}=e_{c}{}^{\mu}\partial_{\mu}$ and where we are ignoring the factor
$(16\pi G_{N}^{(d)})^{-1}$ for the moment in order to get simpler expressions.
A straightforward calculation using the explicit expression of the Einstein
equation Eq.~(\ref{eq:Eb}) gives
\begin{equation}
  \begin{aligned}
(-1)^{d-1} \imath_{k}\mathbf{L}
    & =
      \imath_{k}\star (e^{a}\wedge e^{b}) \wedge R_{ab}
      +(-1)^{d}\star (e^{a}\wedge e^{b}) \wedge \imath_{k} R_{ab}
      -\tfrac{1}{2}\imath_{k}F\wedge \star F
      -\tfrac{1}{2}F\wedge \imath_{k}\star F
    \\
    & \\
    & =
      k^{a}\mathbf{E}_{a} -\imath_{k}F\wedge \star F     
      +(-1)^{d-1}\star (e^{a}\wedge e^{b}) \wedge  \imath_{k} R_{ab}\,.
  \end{aligned}
\end{equation}
The assumption of invariance under the diffeomorphism generated by $k$ implies
the existence of the Maxwell and Lorentz momentum maps $P_{k}$ and
$P_{k}{}^{ab}$ satisfying the momentum map equations
\begin{subequations}
  \begin{align}
    \label{eq:Lorentzmomentummapequation}
    \imath_{k} R^{ab}
    & =
      -\mathcal{D}P_{k}{}^{ab}\,,
    \\
    & \nonumber \\
    \label{eq:Maxwellmomentummapequation}
    \imath_{k}F
    & =
      -dP_{k}\,.
  \end{align}
\end{subequations}
The first equation is always solved by the \textit{Killing bivector}
$\mathcal{D}^{a}k^{b}$.  Using these equations and integrating by parts, we
find
\begin{equation}
  \begin{aligned}
(-1)^{d-1} \imath_{k}\mathbf{L}
    & =
      k^{a}\mathbf{E}_{a} +dP_{k}\wedge \star F     
      +(-1)^{d-1}\star (e^{a}\wedge e^{b}) \wedge  \mathcal{D}P_{k\, ab}
    \\
    & \nonumber \\
    & =
      d\left[
      -\star (e^{a}\wedge e^{b}) P_{k\, ab}
      +P_{k}\star F\right] +P_{k}\mathbf{E}+ k^{a}\mathbf{E}_{a}\,,
        \end{aligned}
\end{equation}
so that
\begin{equation}
  \mathcal{O}_{k}\imath_{k}\mathbf{L}
  \doteq
  d(-1)^{d}\left[
      \star (e^{a}\wedge e^{b}) P_{k\, ab}
      -P_{k}\star F\right]
    =
\left(\mathcal{O}_{k}\mathbf{Q}[k]\right) \,.
\end{equation}
On the other hand, taking the trace of the Einstein equation Eq.~(\ref{eq:Eb})
\begin{equation}
  \begin{aligned}
    e^{a}\wedge \mathbf{E}_{a}
    & =
    (d-2)\star (e^{c}\wedge e^{d})\wedge R_{cd}
      -\frac{(d-4)}{2}F\wedge \star F
    \\
    & \\
    & =
      (d-2)\left[(-1)^{d-1}\mathbf{L}+\tfrac{1}{2}F\wedge \star F
      \right]
       -\frac{(d-4)}{2}F\wedge \star F\,,
    \\
    & \\
    & =
      (d-2)(-1)^{d-1}\mathbf{L}+F\wedge \star F\,,
  \end{aligned}
\end{equation}
so
\begin{equation}
  \mathbf{L}
  =
  \frac{(-1)^{d}}{d-2}F\wedge \star F + \frac{(-1)^{d}}{d-2}e^{a}\wedge\mathbf{E}\,.
\end{equation}
Then,
\begin{equation}
  \begin{aligned}
\imath_{k} \mathcal{O}_{s} \mathbf{L}
  & =
    \imath_{k}\left[  \frac{(-1)^{d}}{d-2}F\wedge \star F \right]
    \\
    & =
      \frac{(-1)^{d}}{d-2}\left[ \imath_{k}F\wedge \star F
      +F\wedge \imath_{k}\star F\right]\,.
  \end{aligned}
\end{equation}
We can use Eq.~(\ref{eq:Maxwellmomentummapequation}) in the first term. In the
second, profiting from the fact that we are working on-shell, we can use the
dual Maxwell momentum-map equation\footnote{Notice that, in $d$ dimensions,
  $\tilde{P}_{k}$ is a $(d-4)$-form.}
\begin{equation}
    \label{eq:dualMaxwellmomentummapequation}
    \imath_{k}\star F
    \doteq
      -d\tilde{P}_{k}\,,  
\end{equation}
and, integrating by parts and using the equations of motion and Bianchi
identities, we have
\begin{equation}
  \begin{aligned}
\imath_{k} \mathcal{O}_{s} \mathbf{L}
  & =
      \frac{(-1)^{d-1}}{d-2}\left[ dP_{k}\wedge \star F
    +F\wedge d\tilde{P}_{k}\right]
    \\
    & \\
    & =
 d \left\{    \frac{(-1)^{d-1}}{d-2}\left[ P_{k}\star F
    +\tilde{P}_{k}F\right] \right\}\,.     
  \end{aligned}
\end{equation}
Now,
\begin{equation}
  [\imath_{k},\mathcal{O}_{s}]\mathbf{L}
  =
  d (-1)^{d-1}\left\{\star (e^{a}\wedge e^{b}) P_{k\, ab}
   -\frac{(d-3)}{d-2} P_{k}\star F
      + \frac{1}{d-2}\tilde{P}_{k}F
\right\}\,,
\end{equation}
which coincides with the result found in Ref.~\cite{Ortin:2022uxa}.

\section*{Acknowledgments}

The work of TO and MZ has been supported in part by the MCI, AEI, FEDER (UE)
grants PID2021-125700NB-C21 (``Gravity, Supergravity and Superstrings''
(GRASS)) and IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S. The work
of MZ has been supported by the fellowship LCF/BQ/DI20/11780035 from ``La
Caixa'' Foundation (ID 100010434). TO wishes to thank M.M.~Fern\'andez for her
permanent support.

\appendix

\end{document}